OpenStudy (anonymous):

find t such that (1,t) lies on the line containing the points (5,5) and (10,8)

OpenStudy (anonymous):

Line is y=3/5 x+2

OpenStudy (anonymous):

t=3/5+2=13/5

OpenStudy (anonymous):

Don't use computers for such silly problems you will never learn how to do it

OpenStudy (anonymous):

how do you do it notsobright? rhetorically speaking, how are you so space?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

my problem was finding the y intercept. my bad, *smart